SophiaH4602:

Help? Marcus spent 10 hours doing his homework last week. This week he spent 11 hours doing homework. He says that he spent 110% more time doing homework this week. Is he correct?

SophiaH4602:

could you explain why he's correct?

SophiaH4602:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @BRUHBRO marcus is my name- \(\color{#0cbb34}{\text{End of Quote}}\) that's great!

dontsaymyname:

He is incorrect because Marcus only spent 10% more time doing homework, Math as followed~ [ 11 - 10 ] / 10 x 100% = 1 /10 x 100% = 10%

darkknight:

its because 11 hours compared to 10 hours. 11 hours is 110% more than 10 hours bc 11/10 = 1.1 and that in percent is 110%

SophiaH4602:

ok thank you guys this was very helpful!

darkknight:

@sophiaH4602 @dontsaymyname is actually wrong

SophiaH4602:

?

darkknight:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dontsaymyname He is incorrect because Marcus only spent 10% more time doing homework, Math as followed~ [ 11 - 10 ] / 10 x 100% = 1 /10 x 100% = 10% \(\color{#0cbb34}{\text{End of Quote}}\) this is incorrect math, the logic here is a little off "[ 11 - 10 ] / 10 x 100% =" you don't do 11-10 can you explain y you did that? when finding percent you divide so y did u subtract here?

darkknight:

actually i might be a bit confused here, hold on a min

dontsaymyname:

Uhm the way I saw it, jumping from 10 to 11 wouldn't be 110% , I was thinking it would be 10% ? but I could be wrong, I'll delete my answer

SophiaH4602:

👁️👄👁️

dontsaymyname:

Wait, I think I'm still correct 0-0

dontsaymyname:

Anybody else going to chime in, or just watch?

darkknight:

this could be up to interpretation, since it is how much MORE homework i agree its 10% MORE not 110% more but I was saying the math you did might be wrong " [ 11 - 10 ] / 10 x 100% = " what is the 11 and the 10 in this???

SophiaH4602:

there hours

dontsaymyname:

The 11 and 10 is from the equation about how many hrs were spent on hw

dontsaymyname:

LMBOOO I think the answer is 10%

SophiaH4602:

It's an old one

darkknight:

ah nvm I see what you did now, yea you did explain it correctly @dontsaymyname my b

dontsaymyname:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight ah nvm I see what you did now, yea you did explain it correctly @dontsaymyname my b \(\color{#0cbb34}{\text{End of Quote}}\) You successfully confused everyone :> jk

SophiaH4602:

PLEASE STOP

dontsaymyname:

Uhm, idk how this became a side topic post, but the answer is 10% @sophiaH4602

SophiaH4602:

ok thank you

SophiaH4602:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @caydenisdead NUUU ITS 69 \(\color{#0cbb34}{\text{End of Quote}}\) please leave

MiraAngel:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dontsaymyname Uhm, idk how this became a side topic post, but the answer is 10% @sophiaH4602 \(\color{#0cbb34}{\text{End of Quote}}\) ^^^^^ Great job. :)

dontsaymyname:

Aww :)

SophiaH4602:

you can go now ty!